OpenStudy (anonymous):

Consider the functions f(x) = 2x2 + 1 and g(x) = x + 2. What is the value of f[g(3)]?

OpenStudy (anonymous):

\[g(3)=3+2=5\] \[f(5)=2\times 5^2+1=51\]

OpenStudy (anonymous):

Assuming you mean f(x) = 2x^2 + 1 and g(x) = x + 2 g(3) = 3 + 2 = 5 f(g(3)) = f(5) = 2*(5^2) + 1 = 51 Same as satellite73 got so I'm guessing I got it right :)